During the annual black friday sale, the OLX sold a pair of ski boots regularly priced at $245.00, at a discount of 40% . The boots cost $96.00 and expenses are 26% of regular selling price. For how much were the ski boots sold?


1
Expert's answer
2021-03-08T19:04:15-0500

Regular Price of boots =$245

Discount rate =40%


Costs of boots=$96


Total expenses=26% of regulat selling prices="\dfrac{26}{100}\times 245=63.70"


Selling price of Boots = "245-\dfrac{40}{100}(245)=245-98=147"


Price of boots= "147+63.70=210"


Hence sold price of boots is $210.